A 6 m simply supported beam carries an ultimate live load of 50kN at midspan. Determine the plastic moment of the beam.


1
Expert's answer
2021-06-02T07:43:13-0400





External deformation = Internal deformation.

EWD= IWD


EWD= Ultimate load "\times" "\delta"


while "\delta = \frac{l}{2} \theta"


EWD = "w_u \times \frac{l}{2} \theta"


from the diagram

IWD = "M_P\theta + M_P2\theta + M_P0 = 4M_P\theta"


IF IWD = EWD

then "4M_P\theta= w_u \times \frac{l}{2} \theta"


"M_P= \frac{wuL}{8} = \frac{50 \times 6}{8} = 37.5 KNm"